OpenStudy (anonymous):

The base of a right prism is a rhombus with a side of 10 cm and an altitude of 9.6 cm. Find the diagonals of the prism if the altitude of the prism is 12 cm.

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

It is an abomination!

OpenStudy (anonymous):

Read the given problem and observe what are the dimension given? If length, base, height and sides are given the prism is triangular If length, width and height are given the given prism is rectangular. Step 2 : If the triangular prism is given find the dimensions using formulas Area of base A = 12 × l × b Perimeter of base P = S1 + S2 + S3 Surface area of prism = lb + (S1 + S2 + S3)h Volume of prism = 12 × l × w × h Where l = altitude or length, b = base and h = height If the rectangular prism find the dimensions using formulas Area of base = l × w Perimeter of base P = (2l + 2w) Surface area of prism = 2(lw) + (2l + 2w)h volume of prism = l × w × h Diagonal of prism = l2+w2+h2−−−−−−−−−−√ Where l = length, w = width and h = height. Put in the values in the above formulas and get the answer.

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):